kaydademon:

For a polynomial p(x), the value of p(3) is −2. Which of the following must be true about p(x)? A) x−5 is a factor of p(x). B) x−2 is a factor of p(x). C) x+2 is a factor of p(x). D) The remainder when p(x) is divided by x−3 is −2. @xxemilianaxx helpp

AnimeLOL:

i can help you!

AnimeLOL:

i know the awnser

kaydademon:

okay

AnimeLOL:

it D.

AnimeLOL:

:)

AnimeLOL:

here hold on

AnimeLOL:

click on the little white thing across from the blue words!

kaydademon:

But i need the math

AnimeLOL:

xxemilianaxx:

Un momento kylie 😂 just need 1 more equation and im done

AnimeLOL:

bye now!

xxemilianaxx:

If the polynomial p(x) is divided by a polynomial of the form x+k (which accounts for all of the possible answer choices in this question), the result can be written as \[p(x) x+k =q(x)+ r x+k\] where q(x) is a polynomial and r is the remainder. Since x+k is a degree-1 polynomial (meaning it only includes x1 and no higher exponents), the remainder is a real number. Therefore, p(x) can be rewritten as p(x)=(x+k)q(x)+r, where r is a real number. The question states that p(3)=−2, so it must be true that −2=p(3)=(3+k)q(3)+r Now we can plug in all the possible answers. If the answer is A, B, or C, r will be 0, while if the answer is D, r will be −2. A. −2=p(3)=(3+(−5))q(3)+0 −2=(3−5)q(3) −2=(−2)q(3) This could be true, but only if q(3)=1 B. −2=p(3)=(3+(−2))q(3)+0 −2=(3−2)q(3) −2=(−1)q(3) This would be right but only if q(3)=2 C. −2=p(3)=(3+2)q(3)+0 −2=(5)q(3) This could be true, but only if q(3)= −2 5 D. −2=p(3)=(3+(−3))q(3)+(−2) −2=(3−3)q(3)+(−2) −2=(0)q(3)+(−2) This will always be true no matter what q(3) is The final answer is D.

AnimeLOL:

yeah

kaydademon:

Okayy thx

AnimeLOL:

good job!!!!!